OpenStudy (anonymous):

What is the Greatest Common Factor of x4 and x3? x x2 x3 x4

OpenStudy (anonymous):

hint \(x^4=x^3\times x\) and \(x^3=x^3\times 1\)

OpenStudy (anonymous):

3

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

so that is the answer right